OpenStudy (theoneandonly):

What is the equation of the line that has a slope of 3 and goes through point (4, 7)?

OpenStudy (jiteshmeghwal9):

equation form when the line passes through \((x_1,y_1)\) & has the slope m is \[(y-y_1)=m(x-x_1)\]

MARC (marc_d):

u can also use this method : slope intercept form\[y=mx+c\]where m=slope c=y-intercept

MARC (marc_d):

substitute (x,y)=(4,7) and m=3 into the eqn to find the value of c

OpenStudy (theoneandonly):

so the equation could be 7=3*4+c

MARC (marc_d):

@jiteshmeghwal9 method is a bit faster..u can use the slope intercept form to double check ur answer..

MARC (marc_d):

yep,that's right.. @Theoneandonly

MARC (marc_d):

solve it to find the c.. 7=12+c 7-12=12-12+c c=-5

MARC (marc_d):

then form the eqn by substituting m=3 n c=-5 into y=mx+c

OpenStudy (theoneandonly):

so the final answer is 7=3*4+(-5)

MARC (marc_d):

u do not need to put the value for x and y,when u want to form the eqn..

MARC (marc_d):

for example,\[y=3x-7\]

OpenStudy (theoneandonly):

oh ok

MARC (marc_d):

okay..back to ur question.. now u hv the value of m=3 and c=-5.. substitute those values into y=mx+c to form the eqn

MARC (marc_d):

\[y=(3)x+(-5)\]

MARC (marc_d):

Therefore,ur eqn is y=3x-5

OpenStudy (theoneandonly):

ok thank you! So my final answer is y=3x-5

MARC (marc_d):

yeap,that's right! ^

MARC (marc_d):

yw! (: